Consider a home mortgage of

​$150,000

at a fixed APR of

6%

for

30

years.

a. Calculate the monthly payment.

b. Determine the total amount paid over the term of the loan.

c. Of the total amount​ paid, what percentage is paid toward the principal and what percentage is paid for interest.
Answer #1

a.

Loan Amount = $150,000

Time Period = 30 years

Interest Rate = 6%

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 150,000, FV = 0, T = 360, I = 0.06/12]

Monthly Payment = $899.33

b.

Total Amount Paid = 360(899.33) = $323,758.80

c.

Principal Paid = 150,000/323,758.80 = 46.33%

Interest Paid = 1 - 0.4633 = 53.67%